# Functoriality of Odd and Generalized Khovanov Homology in $\mathbb{R}^{3}\times I$

### v1

@@ 3.5   Dependence on Sign Assignments and Crossing Orientations
where $\sigma_{L,o}$ denotes the cellular 2-cochain on $Q$ defined by the $\sigma_{i,j}$. To prove the lemma, we now note that any two solutions $\epsilon$ and $\epsilon^{\prime}$ of (32) must differ by a cocycle and hence, since $Q$ is contractible, by a coboundary $\delta\eta$ of a cellular 00-cochain $𝕜\eta\in C^{0}(Q;\Bbbk^{\times})$. The desired chain isomorphism between the generalized Khovanov brackets of $(L,o,\epsilon)$ and $(L,o,\epsilon^{\prime})$ is now given by the cubical chain map with components $\eta(\alpha)\operatorname{id}_{L_{\alpha}}\colon\llbracket L\rrbracket_{\alpha}\rightarrow\llbracket L\rrbracket_{\alpha}$ (where the term cubical means that the chain map has no nonzero components between different resolutions). ∎
@@ 3.6   Generalized Khovanov homology as a diagram and as a colimit
We can obtain a more concrete link invariant by fixing a bidegree $(i,j)\in\mathbb{Z}^{2}$ and replacing each generalized Khovanov bracket $\llbracket D,o,\epsilon\rrbracket$ that appears in the above construction by the corresponding generalized Khovanov homology group in bidegree $(i,j)$. After modding out by the multiplicative action of $𝕜\Bbbk^{\times}$, these generalized Khovanov homology groups become sets $S(D,o,\epsilon,i,j)$, and thus the invariant $\llbracket L\rrbracket_{N}$ becomes a diagram in the category of sets. Taking the colimit of this diagram now yields a set
$S(L,i,j):=\left.\left(\quad\,\,\,\,\bigsqcup_{\mathclap{(\phi,o,\epsilon)\in N(L)}}S(D,o,\epsilon,i,j)\right)\middle/\left(x\sim g_{\phi,o,\epsilon}^{\phi^{\prime},o^{\prime},\epsilon^{\prime}}(x)\right)\right.$
where $g_{\phi,o,\epsilon}^{\phi^{\prime},o^{\prime},\epsilon^{\prime}}$ denotes the map $S(D,o,\epsilon,i,j)\rightarrow S(D^{\prime},o^{\prime},\epsilon^{\prime},i,j)$ that corresponds to the isomorphism from (35). The set $S(L,i,j)$ can be defined for arbitrary links in $\mathbb{R}^{3}$, but not for arbitrary links in $S^{3}$. By the results of this paper, it is functorial in the sense that any smooth link cobordism $F\colon L\rightarrow L^{\prime}$ in $\mathbb{R}^{3}\times[0,1]$ induces a map $S(L,i,j)\rightarrow S(L^{\prime},i,j+\chi(F))$.
@@ 3.7   Invariance under Reidemeister moves
Lemmas 5 through 7 below were shown by Putyra [Pu2015].
@@
If in the above diagrams the orientation of the crossing is flipped, then the orientations of the saddles in the differentials will be flipped as well, but the orientations of the saddles in the chain maps would still the be the ones dictated by Convention 2.
In diagram (36), $S(L_{0},\alpha)$ is the integer
(38) $S(L_{0},\alpha):=\frac{c(L_{0},\alpha)+c(L_{0},\zeta)+deg(\alpha)}{2}$
where $L_{0}$ denotes the link diagram in the bottom row, and $c(L_{0},\alpha)$ and $c(L_{0},\zeta)$ denote respectively the numbers of circles in the resolution $\alpha$ and in the all-zero resolution $\zeta$.
The function $S(L_{0},-)$ appeared in similar form in [Ma2014]. It is defined such that across a merge differential in the resolution cube of $L_{0}$, the increase in the degree and the decrease in the number of components cancel out, while across a split differential, they add up to increment $S(L_{0},\alpha)$ by $1$. The term $c(L_{0},\zeta)$ is to ensure that $S(L_{0},\alpha)$ is an integer. Note that the choice of this term is somewhat arbitrary. We could instead replace it by its negative, or replace $\zeta$ by any fixed resolution of $L_{0}$, such as the oriented resolution. This would at most change the chain maps in (36) by an overall factor of $\pi$.
@@ 3.8   Equivalence of type X and type Y assignments
The see that the same is true for components coming from homotopies, we note that the relevant homotopies are given by the death in $f$ accompanied by the sign $-a\pi\epsilon_{0,0\star\alpha}$, and by the birth in $g$, accompanied by the sign $-\epsilon_{0,1\star\alpha}$, where we are referring to the notations from (39). The same analysis that we used for the maps $f$ and $g$ thus Shows that the components coming from these homotopies Are also the same in $\mathcal{R}(\Phi^{Y}_{F^{\prime}})$ and $\Phi^{X}_{F}$.
If we could show that $\mathcal{R}$ is naturally isomorphic to the identity functor, then Theorem 2 would follow from equations (48) and (50) and from the preceding lemmas. We will instead show a weaker property of $\mathcal{R}$.
@@
The diagram $C$ in $\mathbb{R}^{3}$ consists of nested circles. We can enumerate the circles in $C$, and divide the circles into disjoint sets corresponding to which other circle they are directly nested inside. The diagram $\mathcal{R}(C)$ can inherit the enumeration from $C$ and thus for each set of circles $B_{i}$ with $k_{i}$ circles, the cobordism $S_{C}$ associates an element of $\mathcal{B}_{k_{i}}$ the braid group with $k_{i}$ generators. Overall to $C$, $S_{C}$ associates an element of the group $\prod_{i\in I}\mathcal{B}_{k_{i}}$. If all the circle in a set $B_{i}$ have no circles nested inside them, the four-tube relation gives us the following:
@@
In the braid group this corresponds with the extra relation that generators square to identity. Making this additional identification we are left with $\mathcal{S}_{k_{i}}$ the symmetric group with $k_{i}$ generators. If we could do this for each $B_{i}$, $S_{C}$ would induce an element of $\prod_{i\in I}\mathcal{S}_{k_{i}}$ in turn indicating that all choices for $S_{C}$ are equivalent as they correspond to the identity permutation of the circles.
@@
In the event that two circles in the set $B_{i}$ themselves contain circles consider the following decomposition
(52) [Uncaptioned image]
In essence working from the inside out, the cobordisms inside of a cobordism can be “absorbed” into the wall of the outer cobordism we can then leave the “B” portion of the decomposition below and push the “A” portion above the twisted cobordisms. After we apply the earlier relation, we can slide all of the “B” and “A” sections back together without incurring any signs as we never need to change the orientation on a saddle or change the chronology. ∎
We cannot build a chain map between $\llbracket D^{\prime},o^{\prime},\epsilon^{\prime}\rrbracket$ and $\mathcal{R}(\llbracket D^{\prime},o^{\prime},\epsilon^{\prime}\rrbracket$ by just using the map $S_{C}$ on each resolved crossing diagram $C$ as it would not commute with the differential. If we make appropriate choices of $S_{C}$ it is clear that a merge cobordism will commute with the reflection map, while a split cobordism induces a $\pi$ associated with the reflection reversing the orientation on the split. It follows that there is a cubical chain isomorphism
(53) $\Psi\colon\llbracket D^{\prime},o^{\prime},\epsilon^{\prime}\rrbracket\longrightarrow\mathcal{R}(\llbracket D^{\prime},o^{\prime},\epsilon^{\prime}\rrbracket)$
with components $\Psi_{\alpha}:=\pi^{S(D^{\prime},\alpha)}S_{D^{\prime}_{\alpha}}$ where the exponent $S(D^{\prime},\alpha)\in\mathbb{Z}$ is defined as in (38).
@@
The chain isomorphism in (53) is natural, in the sense that it intertwines the maps $\Phi^{Y}_{F^{\prime}}$ and $\mathcal{R}(\Phi^{Y}_{F^{\prime}})$ up to an overall invertible scalar.
@@
We leave the proof of the lemma to the reader as it is simlar to the computations in Subsection 4.4. It is sufficient to verify that for $F$ an elementary link cobordism $\Phi^{Y}_{F^{\prime}}\stackrel{{\scriptstyle\pi}}{{=}}\Psi_{\alpha}^{-1}\circ\mathcal{R}(\Phi^{Y}_{F^{\prime}})\circ\Psi_{\alpha}$. It is helpful to note that the process of reflecting a death cobordism induces multiplication by $\pi$ as it reverses the orientation on the death. ∎
@@
between the type Y complex $\llbracket D,o,\epsilon\rrbracket$ and the type X complex $\llbracket D,o,\epsilon^{\prime}\rrbracket$. Moreover, Lemmas 8, 9, and 11 imply that this homotopy equivalence is natural, which proves Theorem 2 As an immediate corollary, we obtain a correct proof of the following lemma, which was first stated as Lemma 2.4 in [ORS2007].
@@ 3.9.4   Reidemeister Type Cobordisms
4   The Functoriality of Generalized Khovanov Homology up to Sign
We now have all the elements in place to prove the following theorem.
@@ 4   The Functoriality of Generalized Khovanov Homology up to Sign
Odd Khovanov homology extends to a functor from the category $\mathcal{C}ob^{4}$ to the category $\textnormal{Kom}^{b}_{\pm,\pi}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$.
An immediate consequence of Theorem 5 is the main result of the paper, which is realized by setting $\pi=-1$ in the generalized Khovanov Bracket.
Corollary 2.
Odd Khovanov homology extends to a functor from the category $\mathcal{C}ob^{4}$ to the category $\textnormal{Kom}^{b}_{\pm}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{odd}))$.
In Section 3 we specified what generalized Khovanov homology assigns to movies of link cobordisms. The entirety of this section is devoted to the proof of Theorem 5, where we show how our construction respects all the possible basic ambient isotopies of $\mathcal{C}ob^{4}$. We will begin with a discussion of the type I movie moves, followed by a general argument to show that each side of the type II movie moves produces homotopic chain maps. Next, we will examine type III movie moves with individual arguments in the forward direction, in the reverse direction, and for alternative variants. Finally, we will show the functoriality of odd Khovanov homology with respect to chronological movie moves. We do not need to spend additional time ensuring our functor respects equivalences of planar ambient isotopy cobordisms. Such equivalences are equalities in both $\mathcal{C}ob^{4}$ and $\mathcal{C}ob^{3}$, and are thus respected by our functor.
@@ 4.1   Type I Movie Moves
Figure 4.1. Movie moves 1 through 5
The left-hand sides of the first five movie moves are given by doing and undoing a Reidemeister move, while the right-hand sides (which are not shown in Figure 4.1) are given by trivial movies of identity cobordisms. Since the chain maps assigned to Reidemeister moves are homotopy equivalences, it is clear that the left-hand sides induce chain maps homotopic to the identity, and thus our functor respects these movie moves.

### v2

@@ 3.5   Dependence on Sign Assignments and Crossing Orientations
where $\sigma_{L,o}$ denotes the cellular 2-cochain on $Q$ defined by the $\sigma_{\ldots{\star}\dots\star\ldots}$. To prove the lemma, we now note that any two solutions $\epsilon$ and $\epsilon^{\prime}$ of (34) must differ by a cocycle and hence, since $Q$ is contractible, by a coboundary $\delta\eta$ of a cellular 00-cochain $𝕜\eta\in C^{0}(Q;\Bbbk^{\times})$. The desired chain isomorphism between the generalized Khovanov brackets of $(L,o,\epsilon)$ and $(L,o,\epsilon^{\prime})$ is now given by the cubical chain map with components $\eta(\alpha)\operatorname{id}_{L_{\alpha}}\colon\llbracket L\rrbracket_{\alpha}\rightarrow\llbracket L\rrbracket_{\alpha}$ (where the term cubical means that the chain map has no nonzero components between different resolutions). ∎
@@ 3.6   Generalized Khovanov homology as a diagram and as a colimit
We can obtain a more concrete link invariant by fixing a bidegree $(i,j)\in\mathbb{Z}^{2}$ and replacing each generalized Khovanov bracket $\llbracket D,o,\epsilon\rrbracket$ that appears in the above construction by the corresponding generalized Khovanov homology group in bidegree $(i,j)$ (here $i$ denotes the cohomological grading and $j$ denotes the quantum grading). After modding out by the multiplicative action of $𝕜\Bbbk^{\times}$, these generalized Khovanov homology groups become sets $\mathcal{S}(D,o,\epsilon,i,j)$, and thus the invariant $\llbracket L\rrbracket_{N}$ becomes a diagram in the category of sets. Taking the colimit of this diagram now yields a set
$\mathcal{S}(L,i,j):=\left.\left(\quad\,\,\,\,\bigsqcup_{\mathclap{(\phi,o,\epsilon)\in N(L)}}\mathcal{S}(D,o,\epsilon,i,j)\right)\middle/\left(x\sim g_{\phi,o,\epsilon}^{\phi^{\prime},o^{\prime},\epsilon^{\prime}}(x)\right)\right.$
where $g_{\phi,o,\epsilon}^{\phi^{\prime},o^{\prime},\epsilon^{\prime}}$ denotes the map $\mathcal{S}(D,o,\epsilon,i,j)\rightarrow\mathcal{S}(D^{\prime},o^{\prime},\epsilon^{\prime},i,j)$ that corresponds to the isomorphism from (37). The set $\mathcal{S}(L,i,j)$ can be defined for arbitrary links in $\mathbb{R}^{3}$, but not for arbitrary links in $S^{3}$. By the results of this paper, it is functorial in the sense that any smooth link cobordism $F\colon L\rightarrow L^{\prime}$ in $\mathbb{R}^{3}\times[0,1]$ induces a map $\mathcal{S}(L,i,j)\rightarrow\mathcal{S}(L^{\prime},i,j+\chi(F))$.
@@ 3.7   Invariance under Reidemeister moves
Lemmas 6 through 8 below were shown by Putyra [Pu2015].
@@
In (38), the quantity $S(L_{0},\alpha)$ is defined as in (32). If in the diagrams in (38) and (39) the orientation of the crossing is flipped, then the orientations of the saddles in the differentials will be flipped as well, but the orientations of the saddles in the chain maps would still the be the ones dictated by Convention 2.
@@ 3.8   Equivalence of type X and type Y assignments
The see that the same is true for components coming from homotopies, we note that the relevant homotopies are given by the death in $f$ accompanied by the sign $-a\pi\epsilon_{0,0\star\alpha}$, and by the birth in $g$, accompanied by the sign $-\epsilon_{0,1\star\alpha}$, where we are referring to the notations from (40). The same analysis that we used for the maps $f$ and $g$ thus Shows that the components coming from these homotopies are also the same in $\mathcal{R}(\Phi^{Y}_{F^{\prime}})$ and $\Phi^{X}_{F}$. ∎
Working in the refined setting from Subsection 3.2, we will now show that the functor $\mathcal{R}$ is related to the identity functor via a $\pi$-supernatural isomorphism. Theorem 2 will then follow from equations (49) and (51) and from the preceding lemmas.
@@
The diagram $C$ in $\mathbb{R}^{3}$ consists of nested circles. We can enumerate the circles in $C$, and divide the circles into disjoint sets corresponding to which other circle they are directly nested inside. The diagram $\mathcal{R}(C)$ can inherit the enumeration from $C$ and thus for each set of circles $X_{i}$ with $k_{i}$ circles, the cobordism $S_{C}$ associates an element of $\mathcal{B}_{k_{i}}$ the braid group with $k_{i}$ generators. Overall to $C$, $S_{C}$ associates an element of the group $\prod_{i\in I}\mathcal{B}_{k_{i}}$. If all the circle in a set $X_{i}$ have no circles nested inside them, the four-tube relation gives us the following:
@@
In the braid group this corresponds with the extra relation that generators square to identity. Making this additional identification we are left with $\mathcal{S}_{k_{i}}$ the symmetric group with $k_{i}$ generators. If we could do this for each $X_{i}$, $S_{C}$ would induce an element of $\prod_{i\in I}\mathcal{S}_{k_{i}}$ in turn indicating that all choices for $S_{C}$ are equivalent as they correspond to the identity permutation of the circles.
@@
In the event that two circles in the set $X_{i}$ themselves contain circles consider the following decomposition
Using this decomposition the cobordisms inside of a cobordism can be “absorbed” into the wall of the outer cobordism leaving us with a sum of cobordism each of which can be seperated into a $B_{i}$ portion below and an $A_{i}$ portion above. For each term of the sum we can leave the $B_{i}$ portion in place, and push the $A_{i}$ up past the braiding of the outside circle incuring an invertible scalar $k_{i}$. We can now apply the earlier relation, then slide the $B_{i}$ and $A_{i}$ sections back together again incurring the same invertible scalar $k_{i}$ as we never need to change the orientations on a saddle or change the chronology when applying the relation. ∎
Refer to caption              Refer to caption
Figure 8. The conjugation of reflected split and death cobordisms with the map $S_{C}$ induce multiplication by $\pi$ relative to the unreflected split and death cobordisms. The vertical grey line denotes the plane of reflection
By making appropriate choices for the cobordisms $S_{C}$, one can check that the $S_{C}$ intertwines births and merges with their reflections, and intertwine deaths and splits with their reflections up to a factor of $\pi$ (see Figure (8)). As such, the $S_{C}$ define a $\pi$-supernatural isomorphism between the identity functor and $\mathcal{R}$. We can extend this $\pi$-supernatural isomorphism to a $\pi$-supernatural isomorphism $\eta$ defined on the refinement of $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ by setting $\eta_{C\{m\}}:=\pi^{m}S_{C}\colon C\{m\}\rightarrow\mathcal{R}(C)\{m\}$ for $C$ as above and $m\in\mathbb{Z}_{2}$. By the discussion at the end of Subsection 3.2, we then get a chain isomorphism
(54) $\Psi:=\eta_{\llbracket D^{\prime},o^{\prime},\epsilon^{\prime}\rrbracket}\colon\llbracket D^{\prime},o^{\prime},\epsilon^{\prime}\rrbracket\longrightarrow\mathcal{R}(\llbracket D^{\prime},o^{\prime},\epsilon^{\prime}\rrbracket)$
with components $\Psi_{\alpha}=\pi^{S(D^{\prime},\alpha)}S_{D^{\prime}_{\alpha}}$, where here $\llbracket D^{\prime},o^{\prime},\epsilon^{\prime}\rrbracket$ denotes the refined generalized Khovanov bracket.
@@
The chain isomorphism in (54) is natural, in the sense that it intertwines the maps $\Phi^{Y}_{F^{\prime}}$ and $\mathcal{R}(\Phi^{Y}_{F^{\prime}})$ up to a possible factor of $\pi$.
@@
By Remark 7, the map $\Phi^{Y}_{F^{\prime}}$ that the link cobordism $F^{\prime}$ induces on the refined generalized Khovanov bracket is homogeneous with respect to the supergrading. Since $\Psi$ is a component of a $\pi$-supernatural transformation (by its definition in (54)), Lemma 12 now follows from equation (33) in the definition of a $\pi$-supernatural transformation. ∎
@@
between the type Y complex $\llbracket D,o,\epsilon\rrbracket$ and the type X complex $\llbracket D,o,\epsilon^{\prime}\rrbracket$. Moreover, Lemmas 9, 10, and 12 imply that this homotopy equivalence is natural up to invertible scalars, which proves Theorem 2. As an immediate corollary, we obtain a correct proof of the following lemma, which was first stated as Lemma 2.4 in [ORS2007].
@@ 3.9.2   Death Cobordisms
Remark 6.
The chain map $\Phi_{F}$ assigned to a death cobordism has the following categorical interpretation. Let $L_{0}$ and $L_{1}$ be as above, and let $D\subset\mathbb{R}^{2}$ be a disk that does not intersect $L_{1}$ and that contains the unique component $c$ in $L_{0}\setminus L_{1}$. Let $\mathcal{C}$ be the category defined in the same way as $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$, but with the requirement that the objects of $\mathcal{C}$ are contained in $\mathbb{R}^{2}\setminus D$, and the morphisms are contained in $(\mathbb{R}^{2}\setminus D)\times I$. The inclusion $\mathbb{R}^{2}\setminus D\rightarrow\mathbb{R}^{2}$ induces a functor $\mathcal{C}\rightarrow\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$, which is faithful by [NW2024, Cor. 88] but not full. In particular, this functor allows us to identify $\mathcal{C}$ with a subcategory of $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$. Now consider the functor $\mathcal{F}\colon\mathcal{C}\rightarrow\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ given by sending an object of $\mathcal{C}$ to its union with $c$, and a morphism to its union with $c\times I$. For each object $x\in\mathcal{C}$, there is an obvious morphism $\eta_{x}\colon\mathcal{F}(x)\rightarrow x$ given by a death cobordism in $\mathbb{R}^{2}\times I$ that annihilates $c$. Working in the supergraded refinement from Subsection 3.2, we can further define a morphism $\eta_{x\{s\}}\colon\mathcal{F}(x\{s\})\rightarrow x\{s\}$ by $\eta_{x\{s\}}:=\pi^{s}\eta_{x}$, where $x\{s\}$ denotes the object $x$ with supergrading shifted by $s\in\mathbb{Z}_{2}$. Using that deaths have odd superdegree, it is then easy to see that the morphisms $\eta_{x\{s\}}$ provide a $\pi$-supernatural transformation between (the supergraded refinements of) the functor $\mathcal{F}$ and the inclusion-induced functor $\mathcal{C}\rightarrow\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$. As explained in Subsection 3.2, any $\pi$-supernatural transformation extends to chain complexes, and the death map $\Phi_{F}$ can now be identified with the component $\Phi_{F}=\eta_{\llbracket L_{1}\rrbracket}\colon\llbracket L_{0}\rrbracket\rightarrow\llbracket L_{1}\rrbracket$ of the extended $\pi$-supernatural transformation.
@@ 3.9.4   Reidemeister Type Cobordisms
Remark 7.
One can check that the chain map assigned to an oriented link cobordism $F\subset\mathbb{R}^{3}\times I$ is homogeneous with respect to the supergrading if one works with the supergraded refinement of the generalized Khovanov bracket defined in Subsection 3.2. Moreover, the superdegree of this chain map coincides with the superdegree of $F$ itself. Here it is understood that a birth cobordism in $\mathbb{R}^{3}\times I$ has superdegree 00, and a death cobordism in $\mathbb{R}^{3}\times I$ has superdegree $1$. A saddle cobordism in $\mathbb{R}^{3}\times I$ has superdegree 00 if it merges two link components into a single link component, and superdegree $1$ if it splits a link component into two.
4   Proof of Functoriality
We now have all the elements in place to prove Theorem 1 from the introduction. More precisely, we will show:
@@ 4   Proof of Functoriality
The generalized Khovanov bracket extends to a functor from the category $\mathcal{C}ob^{4}$ to the category $\textnormal{Kom}^{b}_{\pm h,\pm\pi h}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$.
As an immediate consequence of Theorem 5, we obtain Corollary 1 from the introduction, which states that odd Khovanov homology is functorial in $\mathbb{R}^{3}\times I$ up to sign.
In Section 3 we specified what the generalized Khovanov bracket assigns to movies of link cobordisms. The entirety of this section is devoted to the proof of Theorem 5, where we show how our construction respects all the possible movies moves from Theorem 4. We will begin with a discussion of the type I movie moves, followed by a general argument to show that each side of the type II movie moves produces homotopic chain maps. Next, we will examine type III movie moves with individual arguments in the forward direction, in the reverse direction, and for alternative variants. Finally, we will show the functoriality of odd Khovanov homology with respect to chronological movie moves. We do not need to spend additional time ensuring our functor respects planar ambient isotopy moves. The chain maps induced by the two sides of such moves have the same matrix entries in $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ since the underlying chronological cobordisms are ambient isotopic in $\mathbb{R}^{2}\times I$ via an ambient isotopy that preserves the chronological structure.
@@ 4.1   Type I Movie Moves
Figure 9. Movie moves 1 through 5
The left-hand sides of the first five movie moves are given by doing and undoing a Reidemeister move, while the right-hand sides (which are not shown in Figure 9) are given by trivial movies of identity cobordisms. Since the chain maps assigned to Reidemeister moves are homotopy equivalences, it is clear that the left-hand sides induce chain maps homotopic to the identity, and thus our functor respects these movie moves.